We have seen various exponent laws for powers with the same base. CREATE and SOLVE a problem that says to 'SIMPLIFY' and that:  

a) includes multiplying powers with the same base.

b) includes dividing powers with the same base .

c) includes a power of a power.

d) includes an negative exponent.

e) includes an exponent of 0.

f) includes a rational exponent.

g) combines three of the above exponent laws. 


1
Expert's answer
2021-02-02T01:23:41-0500

Solution


  • Simplify

1)

  • Question: "\qquad\qquad
\begin{aligned}
a^2\times a^3 \times a^\frac{2}{3}
\end{aligned}"
  • Answer:"\qquad\qquad
\begin{aligned}
&=a^{2+3+\frac{2}{3}}\\
&=a^\frac{17}{3}
\end{aligned}"

2)

  • Q:"\qquad\qquad
\begin{aligned}
\frac{(a+b)^8}{(a+b)^5}
\end{aligned}"
  • A:"\qquad\qquad
\begin{aligned}
&=(a+b)^{(8-5)}\\
&=(a+b)^3
\end{aligned}"

3)

  • Q:"\qquad\qquad
\begin{aligned}
\sqrt{(a+b)^8}
\end{aligned}"
  • A:

"\qquad\qquad
\begin{aligned}
&=\bigg[(a+b)^8\bigg]^{\frac{1}{2}}\\
&=(a+b)^{8\times\frac{1}{2}}\\
&=(a+b)^4
\end{aligned}"

4)

  • Q:"\qquad\qquad
\begin{aligned}
(\frac{a}{p})^{-5}
\end{aligned}"
  • A:

"\qquad\qquad
\begin{aligned}
&=\frac{a^{-5}}{p^{-5}}\\
&=\frac{1}{a^5}\times\frac{p^5}{1}\\
&=\frac{p^5}{a^5}\\
&=(\frac{p}{a})^5
\end{aligned}"

5)

  • Q:"\qquad
\begin{aligned}
(r+p)^5 \div(r+p)^5
\end{aligned}"
  • A:

"\qquad\qquad
\begin{aligned}
&=\frac{(r+p)^5}{(r+p)^5}\\
&=(r+p)^{5-5}\\
&=(r+p)^0\\
&=1
\end{aligned}"

f)

  • Q:"\qquad\qquad
\begin{aligned}
\frac{a^{\frac{2}{3}}}{a^{\frac{3}{4}}}\\

\end{aligned}"
  • A:

"\qquad\qquad
\begin{aligned}
&=a^{(\frac{2}{3}-\frac{3}{4})}\\
&=a^{-\frac{1}{12}}\\
&=\frac{1}{a^\frac{1}{12}}
\end{aligned}"

g)

  • Q:"\qquad
\begin{aligned}
\frac{a^8\times b^{\frac{3}{4}}}{(ab)^3\times c^{-6}}
\end{aligned}"
  • A:

"\qquad
\begin{aligned}
&=\frac{a^8\times b^{\frac{3}{4}}}{a^3\times b^3\times c^{-6}}\\
&=\frac{a^{(8-5)}\times b^{(\frac{3}{4}-3)}}{c^{-6}}\\
&=\frac{a^3\times b^{-\frac{9}{4}}}{c^{-6}}\\
&=\frac{a^3\times c^6}{b^{\frac{9}{4}}}\\
&=\frac{a^3\times (c^2)^3}{b^{\frac{9}{4}}}\\
&=\frac{(ac^2)^3}{b^{\frac{9}{4}}}

\end{aligned}"